Ann has 1.70 in dimes and nickles. She has two more dimes than nickles. How many nickles does she have.
Answer by stanbon(75887)

Ann has 1.70 in dimes and nickles. She has two more dimes than nickles. How many nickles does she have.
---------------------------------
Value Equation: 10d + 5n = 170 cents
Quantity Eq...: d = n + 2
----------------------------------
Substitute to solve for "n":
10(n+2) + 5n = 170
10n + 20 + 5n = 170
15n = 150
n = 10 (# of nickels)
---
Substitute into d = n+2 to solve for "d":
d = 10+2 = 12 (# of dimes)
===============================
Cheers,
Stan H.